Determine a formula for the arithmetic sequence that begins 1.1, 3.3, 5.5. What is the 10th term of this sequence?
I literally added the 2.2 until I got to the 10th term which is 20.9, but I have no idea on how to construct a real formula. Any help would be greatly appreciated.(:

Answer by user_dude2008(1862)

Formula is 2.2n+1.1 and n starts at 0

1st term: n=0 ----> 2.2(0)+1.1=0+1.1=1.1

2nd term: n=1 ----> 2.2(1)+1.1=2.2+1.1=3.3

3rd term: n=2 ----> 2.2(2)+1.1=4.4+1.1=5.5

...

10th term: n=9 ----> 2.2(9)+1.1=19.8+1.1=20.9


Answer: 10th term is 20.9